OpenStudy (anonymous):

how do u find the length and width of .. x^2+5x+4

OpenStudy (lgbasallote):

let me guess...x^2 + 5x + 4 is the area of a rectangle?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

factor out x^2 + 5x + 4 what do you get?

OpenStudy (anonymous):

how do u factor out? is it (x+4)(x+1) ?

OpenStudy (lgbasallote):

lol yeah..i miswrote the problem but that's the factored form

OpenStudy (lgbasallote):

therefore..those are your length and width! simple isnt it ;)

OpenStudy (anonymous):

i put that but this is what my teacher said "You factored. You need to answer the question. What would be the length and width?"

OpenStudy (lgbasallote):

length = x+4 width = x+1 it can also be: length = x+1 width = x+4 unless there were other specifications?

OpenStudy (anonymous):

i think thats right.. imma try it out.